OpenStudy (mysticfury):

Complete the equation for the linear function whose graph contains the point (−1, 2) and has a slope of 0.9. y =____x + ____ i imagine it would start like y = 0.9x + ____ maybe y = 0.9x + 2 ??

OpenStudy (anonymous):

You are very close! The slope is as you say 0.9. It means that if you increase x by 1, you increase y by 0.9. However, if y = 0.9x + 2, then what happens if you plug in x = -1. You do not get 2, but rather y = 0.9*(-1) + 2 = 1.1.

OpenStudy (anonymous):

Therefore, + 2 cannot be right. You should have gotten 2, since the line was supposed to go through (-1,2)

OpenStudy (mysticfury):

that makes sense, but then how do i figure out what goes in that space?

OpenStudy (mysticfury):

what number*

OpenStudy (anonymous):

Ok, good. Well, what if we say that the right number is a, so that y = 0.9*x + a You know that y = 2 at x = -1

OpenStudy (anonymous):

If you solve for a, then you get the constant term :)

OpenStudy (mysticfury):

okay let me do that, and see what i get

OpenStudy (campbell_st):

the slope intercept form of the line is \[y = mx + b\] m = slope which you know and b = y- intercept so your line is \[y = 0.9x + b\] to find the value of b, use the point x = -1 and y = 2 substitute them and solve for b

OpenStudy (anonymous):

Could not agree more, campbell_st :)

OpenStudy (mysticfury):

hmm okay i got 2 = 0.9 x -1 + 2.9

OpenStudy (campbell_st):

correct

OpenStudy (campbell_st):

so now you know b... and have the equation of the line

OpenStudy (mysticfury):

can you help with another one @SfunnSneffanS ?

OpenStudy (anonymous):

Sure :)